Determine the equation of an exponential function of form y=ab^x that passes through the points (2, 4.5) and (5, 121.5)
Answer by Boreal(15235)

4.5=ab^2
121.5=ab^5
a=4.5/b^2=121.5/b^5
4.5=121.5/b^3
b^3=121.5/4.5
b=3
a=0.5
works in the first
121.5=0.5*3^5, and that works, too.
y=0.5*3^x